OpenStudy (dls):

How would you add 2+2?

OpenStudy (dls):

I am guessing 4

OpenStudy (dls):

@yrelhan4

OpenStudy (dls):

someone please check

OpenStudy (dls):

I used reverse integration on this,and suprisingly it works

OpenStudy (anonymous):

(1+1)+(1+1)=((2)+1)+1=(3)+1=4

OpenStudy (dls):

not clear

OpenStudy (mayankdevnani):